OpenStudy (anonymous):

integral (1/(pix -4)^2 ) I want to know how to answer this question?

OpenStudy (anonymous):

what's pix?

OpenStudy (anonymous):

pi times x

OpenStudy (agreene):

\[\int\frac1{(\pi x-4)^2}dx\] Is this correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

sub u=(πx−4) and du= π dx

OpenStudy (agreene):

Let: u = pi * x-4 du = pi dx \[\therefore \pi^{-1}\int u^{-2}du=-\pi*u^{-1}+C=-\frac{1}{\pi u}+C\] plug back in: \[\frac1{4\pi-\pi^2x}+C\]

OpenStudy (anonymous):

i would use substitution u=πx−4 du=π dx so this would be \[ 1/\pi \int\limits 1/ u ^{2} du \] the integral of 1/u^2 is -1/u so you get \[1/4\pi- \pi^2x + C\] Substitute back for u = pi x-4: = 1/(4 pi-pi^2 x)+constant

OpenStudy (anonymous):

I don't understand how do you get that 1/pi before the integral?

OpenStudy (anonymous):

it comes from the substitution